# BrightDreamer: Generic 3D Gaussian Generative Framework for Fast Text-to-3D Synthesis

### v1

@@ 3.4 Optimization
$\nabla_{\theta}\mathcal{L}_{\mathrm{SDS}}(\phi,\mathbf{x}=g_{\theta}(prompt))\triangleq\mathbb{E}_{t,\epsilon}\left[w(t)\left(\hat{\epsilon}_{\phi}\left(\mathbf{z}_{t};y,t\right)-\epsilon\right)\frac{\partial\mathbf{x}}{\partial\theta}\right],$ (8)
where $prompt$ is the input prompt of the generator, $\theta$ is the 3D Gaussians generator’s trainable parameters, $\phi$ is the parameters of denoising network, $x$ is the generated image by process $g_{\theta}(\cdot)$, $w(t)$ is the weighting function with time step $t$ in the denoising schedule, $\epsilon$ is the random noise, $z_{t}=x+\epsilon$, $y$ is the adjusted text according to the sampling view direction, $\hat{\epsilon}_{\phi}(\cdot)$ is the predicted noise.
@@ 4 Experiments
4.1 Implementation Details
All our experiments are conducted on a server with 8 GPUs with 80GB memory. Our codebase is constructed on the PyTorch framework [42] with Automatic Mixed Precision Training (AMP). And we introduce the Gradient Checkpointing technology [9] to save the GPU memory. We use the Adam optimizer [23] to update the parameters of our generator with a constant learning rate of $5\times{10}^{-5}$, $\beta_{1}$ of $0.9$ and $\beta_{2}$ of $0.99$. We train our generator using the images rendered at $512\times 512$ resolution by original rasterization of 3D Gaussian Splatting [22] and use the DeepFloyd IF [51] UNet to calculate the SDS Loss (Eq. 8). The prompt batch size is set to $64$ in total and the camera batch size is set to $4$. We set the freedom $\beta$ (Eq. 4) to $0.2$, and the range of scaling $(a,b)$ (Eq. 7) to $(-9,-3)$. The anchor position is placed as a $64^{3}$ 3D grid, and the resolution of the generated triplane is $256\times 256$.
4.2 Datasets
Vehicles. We construct a vehicle’s prompts set using ChatGPT [^1] https://chat.openai.com/ to generate a description about car details, e.g., "vintage convertible, cherry red, chrome bumpers, white-wall tires". There are 3,026 prompts containing 1,856 unique words in total.
Daily Life. This dataset is constructed by Instant3D [27]. They use the ChatGPT1 to generate more than $17,000$ prompts containing $3,135$ unique words. For example, "A tall man with a beard and wearing a leather jacket is riding a motorcycle".
Animals. This prompt set is introduced by Instant3D [27]. And the structure is "a {species} sitting {item} and wearing {gadget} and wearing a {hat}". Species contains [’wolf’, ’dog’, ’panda’, ’fox’, ’civet’, ’cat’, ’red panda’, ’teddy bear’, ’rabbit’, ’koala’]. Item contains [’in a bathtub’, ’on a stone’, ’on books’, ’on a table’, ’on the lawn’, ’in a basket’, ’null’]. Gadget contains [’a tie’, ’a cape’, ’sunglasses’, ’a scarf’, ’null’]. Hat contains [’beret’, ’beanie’, ’cowboy hat’, ’straw hat’, ’baseball cap’, ’tophat’, ’party hat’, ’sombrero’, ’null’]. There are $3,150$ prompts in total.
4.3 Demonstration of BrightDreamer
To show the quality and view consistency of BrightDreamer, we show some results in Fig. 6. All prompts do not appear in our training set. Additionally, we show much more visual results in our supplementary details.
Table 1: The generation latency (millisecond) and rendering speed (FPS, Frames Per Second) of our generator.
@@ 4.3 Demonstration of BrightDreamer
Figure 7: Vehicle generation comparison. All prompts don’t appear in the training set. The result of Dreamfusion is reproduced by ThreeStudio [^3] https://github.com/threestudio-project/threestudio.
@@ 4.4 Comparison with Other Methods
Inference time. In Tab. 1, we show the inference latency on a single A800 GPU and a single RTX3090 GPU. BrightDreamer can generate a 3D GS representation for any text prompt in less than 0.1 seconds, which shows a large margin improvement, compared to optimization-based methods, which need several hours to optimize for one single prompt. The generated 3D Gaussians can be rendered at over 700 FPS speed.
Figure 8: Visual comparison of daily life. All prompts don’t appear in training set.
Visual comparison results on vehicle prompts. In Fig. 7, we show the comparison results. It’s noteworthy that other methods all need to train different models for different prompts. In contrast, our method can directly infer one-time for a single prompt and all the test prompts do not appear in our training prompt set. Because there is no open-source codebase of Instant3D [27] and ATT3D [33], we only show our result compared to other optimization-based methods. Due to the car generation limitations remaining in Stable Diffusion, we use DeepFloydIF denoising UNet in Eq. 8. However, it can only provide the $64\times 64$ supervision, so there remains some unfairness. We first compare with SoTA NeRF optimization methods, dreamfusion [43] and ProlificDreamer [56]. As you can see, the generation result is extremely unreasonable and chaotic, which shows the difficulty when working on relatively complex prompts. Though DreamGaussian’s training is relatively fast, 5 minutes, the generation result is still relatively bad. Since LucidDreamer [30] and GSGEN [10] introduce the Point-E [40] as their prior or supervision, they can generate the correct shape of a vehicle. However, the LucidDreamer has abnormal light spots, which cause unreal sensing. Our method can deal with arbitrary text prompts in 77 ms, which shows an extremely strong application value. Importantly, you can scale up to check the second two columns in Fig. 7. Looking through the window at the seats inside, there is an irregular phenomenon in the seat of GSGEN and LucidDreamer. But ours looks reasonable, which shows that joint training can make the network have a stronger capability of semantic understanding.
Visual comparison results on daily life prompts. In Fig. 8, we compare our method to Instant3D [27] and other per-prompt optimization methods. As you can see, nearly all per-prompt methods fail to generate reasonable content. Either all entities in the prompt appear, but the interaction relationship is significantly wrong. Either they lack some entities. Though ProlificDreamer [56] and LucidDreamer [30] improve the original SDS loss, which allows them can get more details in the generated content, the consistency to the prompt is damaged a lot. Our BrightDreamer not only uses 77ms to generate but also surpasses other methods in the aspect of rationality far away. Additionally, the generated 3D Gaussians can be rendered over 700 FPS, which is 30 times than other NeRF-based methods, e.g. Instant3D.
Figure 9: The visual comparison of animal generation.
Visual comparison results on animal prompts. In Fig. 9, we compare our method to Instant3D [27] and other per-prompt optimization methods. As you can see, if we only substitute some words in the sentence, our method still can remain the similar style. Specifically, in the first and second two columns, we only change the thing which the teddy bear sitting on. It is hard for per-prompt optimization-based methods to maintain the similarity of other parts. And it is difficult for them to generate all entities that appeared in the input prompts. There are also some wrong understandings of the input prompt, e.g., the rabbit case in GSGEN [10]. In general, our BrightDreamer achieves the best ability of understanding and the highest speed in generation and rendering.
@@ 4.5 Ablation Studies
In Fig. 10, we demonstrate the influence of our network design for training. Compared between Fig. 10 (a) and Fig. 10 (b), our divided triplane generator can reduce the degree of chaos in the space significantly, which shows the necessity of our division. As Fig. 10 (a) and Fig. 10 (c) demonstrate, it is necessary to pass the coordinate into the $F_{shape}$ and $F_{color}$. This design can construct a gradient pathway toward the TSD Network, ensuring more accurate shape formulation.
Figure 10: The key components in our design. All the models are trained to 10, 000 iterations with the same configuration. (a) Our completion design. (b) Replace the three separate generators with a single generator. (c) Don’t input the coordinate to $F_{shape}$ and $F_{color}$.
@@ 5 Conclusion
In this paper, we introduced the first text-driven 3D Gaussians generative framework, BrightDreamer, capable of generating 3D Gaussians within a remarkably low latency of 77ms. To address the challenge of directly generating a vast quantity of Gaussians, millions of them, we innovatively deform anchor positions and use the new positions as the centers of 3D Gaussians, in response to the input prompt. This approach effectively circumvented the obstacle of generating a large number of positions. Regarding network architecture, we conducted a thorough reevaluation of the triplane generation process and introduced an improved alternative strategy. Our largest contribution is poised to significantly advance the field of generalized 3D generation, offering a novel and efficient pathway to creating 3D assets from text prompts immediately. A large number of experiments prove that our BrightDreamer has a strong complex semantic understanding capability and strong generalization ability.
Future work. Both Instant3D and our approach are constrained to generating a singular 3D content for each input, which is different from GAN and diffusion models. These traditional methods offer the flexibility to vary the generation outcomes by altering the input noise. We aim to address this challenge in our future research endeavors, with the goal of introducing variability and diversity for a single input.
Appendix 0.A Training Details of BrightDreamer
For a clearer illustration of our training pipeline, we provide the details in Algorithm. 1. We train 36 hours for the vehicle prompts set, 60 hours for the daily life prompts set, and 30 hours for the animal prompts set on a server with 8 80GB GPUs.
@@ Appendix 0.A Training Details of BrightDreamer
        $cameras$ ${}$ randomly sample $C$ cameras on upper sphere;
        $rendered\_images$ ${}$ splatting($3D\_GS[i]$, $cameras$);
@@
        $loss$ ${}$ $loss$ + SDS($rendered\_images$, $texts\_dir$);   # Eq. 8
@@
Appendix 0.B Inference Procedures of Our BrightDreamer
@@
Appendix 0.C More Visual Results of Ablation Studies
In Fig. 12, We provide more visual results in ablation studies.
Figure 11: More ablation visual results. All models were trained to $10,000$ iterations with the same other configurations. (a) Our completion design. (b) Replace the three separate generators with a single generator. (c) Don’t input the coordinate to $F_{shape}$ and $F_{color}$. The prompt of the first line is "Muscle car, silver moon, hot hatch performance ". The prompt of the second line is "Vintage station wagon, cobalt blue, space-efficient design, economical ".
Appendix 0.D Discussion about Generalization
@@
Appendix 0.E More Visual Results

### v2

@@ 3.4 Optimization
$\nabla_{\theta}\mathcal{L}(\phi,\mathbf{x}=g_{\theta}(y))\triangleq\mathbb{E}_{t,\epsilon}\left[w(t)\left(\hat{\epsilon}_{\phi}\left(\mathbf{z}_{t};y^{\prime},t\right)\!-\epsilon\right)\!\frac{\partial\mathbf{x}}{\partial\theta}\right],$ (6)
where $y$ is the input prompt embedding of the generator, $\theta$ denotes the trainable parameters of 3D Gaussians generator, $\phi$ denotes the parameters of denoising network, $x$ is the generated image by process $g_{\theta}(\cdot)$, $w(t)$ is the weighting function with time step $t$ in the denoising schedule, $\epsilon$ is the random noise, $z_{t}$ is the noisy latents encoded from $x$, $y$ is the input prompt, and $y^{\prime}$ adjusted text according to the sampling view direction, $\hat{\epsilon}_{\phi}(\cdot)$ is the predicted noise.
@@ 4 Experiments
This section first outlines the implementation details, inference speed, parameter count, and training dataset. The following comparison experiments focus on four aspects: (I) comparison with per-prompt text-to-3DGS, (II) fine-tuning the generated 3DGS model and then comparing with per-prompt text-to-3DGS, (III) comparison with amortized text-to-NeRF, (IV) comparison with text-to-image-to-3DGS methods. Finally, we present interesting findings regarding unseen words and conduct ablation studies to validate the network design. It’s worth noting that all prompts in the experiments don’t appear in the training dataset.
4.1 Implementation Details
Our codebase is constructed on the PyTorch framework [52] with Automatic Mixed Precision Training (AMP) and Gradient Checkpointing technology [9]. We use the Adam optimizer [27] with a constant learning rate of $5\times{10}^{-5}$, $\beta_{1}$ of $0.9$ and $\beta_{2}$ of $0.99$. We train our generator using the DeepFloyd IF [64] UNet to calculate the SDS Loss (Eq. 6). The prompt batch size is set to $64$ and the camera batch size is set to $4$. All our experiments are conducted on a server with 8 GPUs with 80GB memory. We set the freedom $\beta$ (Eq. 2) to $0.2$, and the range of scaling $(a,b)$ (Eq. 5) to $(-9,-3)$. The anchor position is placed as a $64^{3}$ 3D grid, and the resolution of the generated triplane is $256\times 256$. The generator is trained on a single prompt set including vehicle, daily life, and animal descriptions. We provide the training pseudo code in supplementary materials.
4.2 Demonstration of BrightDreamer
@@
Table 1: The generation latency (millisecond) and rendering speed (FPS, Frames Per Second).
We show some results in Fig. 6. All prompts don’t appear in the training set. We also show much more results in supplementary material. In Tab. 1, we show the inference latency on a single A800 GPU and a single RTX3090 GPU, which shows a large margin improvement, compared to optimization-based methods, which need several hours. Our generator contains about 500M trainable parameters.
@@ 4.3 Comparison with other methods
We compare our method with previous text-to-3DGS methods. The comparison experiments setting includes four aspects: (I) comparison with per-prompt text-to-3DGS, including DreamGaussian [65], GSGEN [10] and LucidDreamer [37] (II) fine-tuning the generated 3DGS model, and then compare with per-prompt text-to-3DGS, (III) comparison with amortized text-to-NeRF, including Instant3D [33], (IV) comparison with text-to-image-to-3DGS methods, including LGM [66] and GRM [80]. Besides, we present the quantitative results of human preference.
4.3.1 Per-prompt Text-to-3DGS
Figure 7: Comparison with per-prompt optimization methods.
We select the current SoTA open-source text-to-3DGS methods, including DreamGaussian [65], GSGEN [10] and LucidDreamer [37] for comparison. In this setting, we directly input the text prompt to BrightDreamer and use the same prompt to optimize the 3D GS model in other methods. As shown in the first four columns of Fig. 7, the other methods are hard to generate semantic-aligned 3D GS models from complex text prompts, primarily due to their limited capability in comprehending intricate text inputs. A similar conclusion can be found in the works [21, 92, 11, 4, 13]. It is hard to optimize a 3D GS model for a prompt containing multiple objects. Our method, trained across a diverse distribution of multiple text prompts, exhibits a robust capability to comprehend complex prompts effectively.
4.3.2 Quickly Finetuning after generation
We demonstrate that the 3D GS model generated by BrightDreamer can be efficiently fine-tuned using per-prompt optimization methods to enhance quality. Specifically, we use the ISM [37] loss to optimize the generated 3D GS model for several hundred iterations. As evidenced in the last column of Fig. 7, the third and sixth columns of Fig. 8, the model’s quality of details significantly improves within just several minutes.
4.3.3 Amortized Text-to-NeRF
Figure 8: Comparison with amortized text-to-NeRF method.
We compare BrightDreamer with SoTA open-source amortized Text-to-NeRF methods, Instant3D [33]. It is worth noting that our method is the first amortized Text-to-3DGS method. Moreover, our method inherits numerous advantages of 3D GS over NeRF. Since ATT3D [42] and Latte3D [76] don’t open their code and demo, we don’t compare with them. As shown in Fig. 8, our semantic alignment is better. Furthermore, benefiting from the training efficiency of 3D GS, we can quickly fine-tune the generated 3D GS model to enhance the details.
4.3.4 Text-to-image-to-3DGS
Figure 9: Comparison with text-to-image-to-3DGS methods.
Besides the direct generation of text-to-3DGS, alternative methods also exist to achieve this goal. For instance, LGM [66] and GRM [80] first generate the reference images using a text-to-image diffusion model, followed by the creation of a 3D model from these images. However, since the second phase of image-to-3DGS is trained on the limited 3D dataset, it is hard for them to generate the 3D GS models for complex prompt input. As shown in Fig. 9, we demonstrate the comparison result. Furthermore, since these methods require an intermediate image generation step, they are considerably slower than our BrightDreamer approach.
4.3.5 Quantitative Comparison
Method Instant3D [33] LucidDreamer [37] Ours
Preference 37.3% 7.2% 55.4%
Table 2: Human Preference.
As shown in Tab. 2, we provide the percentage numerical comparison of human preference choice. We present participants with five rendered videos along with the corresponding text prompts generated by five baseline models, allowing them to select their preferred option for each case. Most of the users expressed a preference for the content generated by our model.
4.4 Discussion about Generalizability
Figure 10: Demonstration of word-level generalizability. The word “banana” never appears in the training set.
Beyond sentence-level generalizability, our findings indicate that BrightDreamer can also correctly interpret some unseen words. As demonstrated in Fig. 10, although the word ”banana” is absent from our training dataset, BrightDreamer successfully identifies its color. This capability not only suggests a potential for broad word-level generalizability but also confirms that BrightDreamer effectively learns the mapping from text distribution to 3D GS distributions.
@@ 4.5 Ablation Studies
As shown in Fig. 11, we validate our network design for training. Compared between Fig. 11 (a) and Fig. 11 (b), our divided triplane generator can reduce the degree of chaos in the space significantly, which shows the necessity of our division. As Fig. 11 (a) and Fig. 11 (c) demonstrate, it is necessary to pass the coordinate into the $F_{shape}$ and $F_{color}$. This design can construct a gradient pathway toward the TSD Network, ensuring more accurate shape formulation.
Figure 11: Ablation studies. All models are trained for 10,000 iterations (incomplete training) with the same configuration: (a) Our complete design, (b) Single generator replacing three separate generators, (c) Coordinate input removed from $F_{shape}$ and $F_{color}$.
@@ 5 Conclusion
In this paper, we introduce the first text-driven 3D Gaussians generative framework, BrightDreamer, capable of generating 3D Gaussians within a remarkably low latency of 77ms. To efficiently generate millions of 3D Gaussians, we innovatively deform anchor positions and use these new positions as centers based on the input prompt. This approach successfully overcomes the challenge of generating a large number of positions. Regarding network architecture, we thoroughly reevaluate the triplane generation process and introduce an improved alternative strategy. Our key contribution is to greatly enhance the generalized 3D generation, offering a novel and efficient way to create 3D assets from text prompts instantly. Extensive experiments demonstrate that BrightDreamer possesses strong semantic understanding and generalization abilities. Due to the limited 3D data, developing methods based on 2D diffusion is an important direction.
Future work. The spatial resolution of the generated 3D model is relatively low, resulting in the lack of fine-grained detail. Besides, since the LGM [66] and GRM [80] can generate highly detailed single-object 3D GS models while our method can generate semantic-rich 3D GS models, developing a method that combines these two advantages is an interesting research direction.
Appendix A Training Details of BrightDreamer
For a clearer illustration of our training pipeline, we provide the details in Algorithm. 1. We train our BrightDreamer for about 2 days on a server with 8 GPUs with 80GB memory. Actually, a lower batch size and less training time can also result in similar quality.
@@ Appendix A Training Details of BrightDreamer
        $cameras$ ${}$ randomly sample $C$ cameras;
        $images$ ${}$ render($3D\_GS[i]$, $cameras$);
@@
        $loss$ ${}$ $loss$ + SDS($images$, $texts\_dir$);   # Eq. 6
@@
Appendix B Inference Procedures of Our BrightDreamer
@@
Appendix C Discussion about Generalization
@@
Appendix D More Visual Results